OpenStudy (batmaaaanxd):

[Plz help - I'll medal and fan] Solve algebraically: x/2+y/4=3 (1) 2x+y=3/2

OpenStudy (batmaaaanxd):

@jhonyy9 @inkyvoyd

OpenStudy (batmaaaanxd):

@Aveline

Directrix (directrix):

Multiply each term of the first equation by 4. Post what you get, okay? x/2 + y/4 = 3

OpenStudy (batmaaaanxd):

@Directrix Thanks for replying. Oh and I got 2x + y = 12..?

Directrix (directrix):

2x + y = 12 2x + y = 3/2 ----------- The left sides of the two equations are the same. The right sides are not. If you subtract the second from the first, you get: 0 + 0 = 12 - 3/2 which is not true.

Directrix (directrix):

The lines do not have a common point of intersection. Look at the attached graphs.

Directrix (directrix):

@batmaaaanxD What is the relationship between the graphs of the two equations?

OpenStudy (batmaaaanxd):

@Directrix Thanks for going through all this trouble but I just found that the question was wrong. The second equation is 2x-3y=12. I can solve it now, thank you!